### v1

# PARCO: Learning Parallel Autoregressive Policies for Efficient Multi-Agent Combinatorial Optimization

Abstract
Multi-agent combinatorial optimization problems such as routing and scheduling have great practical relevance but present challenges due to their NP-hard combinatorial nature, hard constraints on the number of possible agents, and hard-to-optimize objective functions. This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel approach that learns fast surrogate solvers for multi-agent combinatorial problems with reinforcement learning by employing parallel autoregressive decoding. We propose a model with a Multiple Pointer Mechanism to efficiently decode multiple decisions simultaneously by different agents, enhanced by a Priority-based Conflict Handling scheme. Moreover, we design specialized Communication Layers that enable effective agent collaboration, thus enriching decision-making. We evaluate PARCO in representative multi-agent combinatorial problems in routing and scheduling and demonstrate that our learned solvers offer competitive results against both classical and neural baselines in terms of both solution quality and speed. We make our code openly available at https://github.com/ai4co/parco.
[…]
1 Introduction
Combinatorial optimization (CO) problems, such as routing and scheduling problems, involve determining an optimal sequence of actions in a combinatorial space and have applications ranging from warehouse operatons (Xie et al., 2023) to network design (Chabarek et al., 2008) and safety-critical systems (Girardey et al., 2010). CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time, i.e., they are NP-hard (Jünger et al., 1995). Multi-agent settings gained significant interest due to their applicability in realistic scenarios, such as path finding (Reijnen et al., 2020), drone routing (Ann et al., 2015), disaster management (Bektas, 2006; Cheikhrouhou and Khoufi, 2021) and order delivery (Yakıcı and Karasakal, 2013; Archetti and Bertazzi, 2021), but present even more challenges due to additional constraints and different optimization objectives, including minimizing a global lateness objective or the makespan (Mahmoudinazlou and Kwon, 2024).
While traditional algorithmic methods have significantly contributed to solving a range of problems (Laporte and Osman, 1995; Hejazi and Saghafian, 2005), these approaches often concentrate on single-agent scenarios. With the advent of modern computational techniques, neural network-based approaches have started to yield promising results for complex CO problems in a field known as Neural Combinatorial Optimization (NCO). In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with environments instead of relying on (costly) labeled datasets in the shape of optimal solutions (Bello et al., 2016; Kwon et al., 2020; Kim et al., 2023a).
Among NCO methods, Autoregressive (AR) sequence generation has gained attention for its ability to manage hard constraints (Kool et al., 2018; Kwon et al., 2021). In the context of NCO, this capability is crucial for addressing complex problems with multiple construction constraints, such as the precedence constraints prevalent in scheduling (Zhang et al., 2020a) as well as pickup and delivery problems (Savelsbergh and Sol, 1995; Parragh et al., 2008). However, one issue of autoregressive sequence generation is the high latency associated with it, especially when considering large problem instances. The problem of high inference latency of AR methods is especially prevalent in the domain of large language models, as multiple stacks of transformer layers have to be computed sequentially for each individual token (Bae et al., 2023).
Figure 1: Solution construction with Autoregressive (top) and Parallel Autoregressive policies (bottom).
Motivated by recent studies on LLMs demonstrating parallel decoding can not only tackle the high generation latency problem of next-token predictors but also improve their answer quality (Qi et al., 2020; Ning et al., 2023; Gloeckle et al., 2024) this paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel approach designed to address multi-agent combinatorial problems efficiently. We propose to apply parallel decoding in the NCO domain to increase solution construction efficiency and effectiveness. In contrast to most previous NCO methods, PARCO constructs solutions in parallel for multiple agents, made feasible by a priority-based conflict handler to manage conflicting decisions of different agents. Importantly, we enable the collaborative behavior of agents through specialized inter-agent Communication Layers, essential for finding high-quality solutions to multi-agent combinatorial problems.
Contributions.
We summarize our contributions as follows:
  • We introduce PARCO, a novel approach for tackling multi-agent CO problems by effectively constructing solutions in parallel via a Multiple Pointer Mechanism.
  • We introduce an efficient Priority-based Conflict-Handling scheme to resolve conflict between agents and enhance decision-making.
  • We design a special Communication Layer to help agents communicate and collaborate to obtain a better state-space representation of partial solutions.
  • We show competitive performance against classical solvers and neural baselines in routing and scheduling problems, both in terms of solution quality and computational efficiency.
2 Related Work
Neural Combinatorial Optimization
Recent advancements in NCO have shown promising end-to-end solutions for combinatorial optimization problems, as highlighted by Bengio et al. (2021) and Yang and Whinston (2023). NCO has led to the development of aligned neural architectures (Bi et al., 2022; Jin et al., 2023; Luo et al., 2023; Kim et al., 2023c), hybrid methods with OR solvers (Li et al., 2021b; Kim et al., 2024a; Yan and Wu, 2024; Ye et al., 2024a; Kim et al., 2024b), multi-level solution pipelines (Ma et al., 2023a; Li et al., 2023; Xiao et al., 2023; Ye et al., 2023), alongside improved training algorithms (Kim et al., 2023c; Jiang et al., 2023; Drakulic et al., 2023; Sun and Yang, 2023; Gao et al., 2023; Xiao et al., 2023; Li et al., 2024b; Wang et al., 2024) to enhance the heuristic search. These innovations have expanded NCO’s application across a wide array of problems (Chen et al., 2023; Kim et al., 2023b; Zhou et al., 2023; Ma et al., 2023b; Luttmann and Xie, 2024). However, integrating appropriate inductive biases requires manual tuning of model architectures and training algorithms, presenting challenges such as computationally intensive training, the need for specialized hardware, and issues with interpretability and generalizability (Liu et al., 2023).
Multi-agent Constructive Methods for NCO
While several seminal works as Vinyals et al. (2015); Kool et al. (2018); Kwon et al. (2020) propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes, such as different capacities or starting locations in routing problems. Zhang et al. (2020b); Falkner and Schmidt-Thieme (2020); Li et al. (2022); Liu et al. (2024c) introduce models to construct solutions of different agents simultaneously. However, the above parallel construction is, in fact, sequential in terms of the model and environment stepping itself, thus not benefitting from true model-side parallel decoding and often lacking in agent collaboration. On the other hand, Son et al. (2024); Zheng et al. (2024) propose to model multi-agent min-max routing as sequential decision-making, allowing for switching the agents after a single-agent solution is complete. Most related to PARCO is Zong et al. (2022), who propose a multi-agent model for the min-sum multi-agent PDP with different decoders. However, this approach is not only tailored to a specific CO problem but also a specific number of agents due to both different decoders for each agent and a non-flexible context embedding, which is dependent on the agent number, while PARCO is more flexible and has a more general communication representation. Moreover, while Zong et al. (2022) manages conflicts by randomly giving precedence to one agent, PARCO learns how to prioritize agents with higher probabilities of selecting a specific node.
3 Background
3.1 Markov Decision Processes
CO problems can be framed as Markov Decision Processes (MDPs). In this formulation, the problem is defined by a set of states $\mathcal{S}$, where each state $s_{t}\in\mathcal{S}$ represents the configuration of the problem at time step $t$. At each step, an agent selects an action $a_{t}$ from the action space $\mathcal{A}$ according to a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$, which maps states to actions. The system then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. The agent reaches the terminal state once it has generated a viable solution for the problem instance $\bm{x}$. Typically, a reward $R$ is only obtained in the terminal state and takes the form of the objective function of the respective CO problem.
3.2 Autoregressive Methods for NCO
Given the sequential nature of MDPs, autoregressive (AR) methods pose a natural choice for the policy $\pi$. AR methods construct a viable solution by sequentially generating actions based on the current state and previously selected actions. Formally, the process can be represented as:
$p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}g_{\theta}(a_{t}|a_{t-1},...,a_{0},\bm{h})$ (1)
where $\bm{h}=f_{\theta}(\bm{x})$ is the encoding of problem instance $\bm{x}$ via encoder network $f$, which is used to decode actions autoregressively via decoder $g_{\theta}$. $p_{\theta}$ is a solver that maps $\bm{x}$ to a solution $\bm{a}$. $\bm{a}=(a_{1},...,a_{T})$ represents the (optimal) actions executed in $T$ construction steps, resulting in a feasible solution to CO problems.
Notably, the same AR scheme has been applied to solve multi-agent CO problems, such as vehicle routing problems (Kool et al., 2018; Kwon et al., 2020), in which a single agent tasked with visiting all nodes under specific route length constraints, distinguishing it from a real multi-agent scenario by focusing on optimizing a solitary agent’s path within set limits. To tackle such problems, AR methods sequentially construct multiple routes as if handling a single agent at a time. However, inducing collaborative behavior among agents using the AR method poses a challenge, as it inherently focuses on one agent at a time. This limitation restricts the method’s ability to directly foster interactions among the agents, which is crucial for optimizing collective outcomes in multi-agent scenarios.
3.3 Training AR Models via Reinforcement Learning
In this work, we focus on Reinforcement Learning (RL) solvers for training $p_{\theta}$ as they can be trained without relying on (often hard-to-obtain) labeled solutions.
We can thus cast the optimization problem as follows:
[…]
where $P(\bm{x})$ is problem distribution, $R(\bm{a},\bm{x})$ is reward of $\bm{a}$ given $\bm{x}$. We can solve Eq. 2 by employing various RL methods such as variations of policy gradients (REINFORCE) (Kwon et al., 2020; Kim et al., 2022).
4 Methodology
4.1 Parallel Multi-Agent Environments
We propose to step multiple actions over the environment simultaneously to enhance efficiency in practice, which can reduce the number of steps required compared to single-agent stepping. In the general multi-agent CO setting, agents $k=1,...,m$ are selecting actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{m})$ from a shared action space $\mathcal{A}$ in parallel at a given decoding step $t$. Given the agent actions $\bm{a}_{t}$, the state of the problem instance transitions from $s_{t}$ to $s_{t+1}$ via some transition function $T:\mathcal{S}\times\mathcal{A}_{1}\times...\times\mathcal{A}_{m}\rightarrow\mathcal{S}$, which usually follows deterministic rules in case of non-stochastic CO problems. After reaching the terminal state, the agents receive a shared reward $R(A,\bm{x})$, with $\bm{x}$ the problem instance and $A=(\bm{a}_{1},...,\bm{a}_{T})$ the sequence of agents actions, which is the objective of the respective CO problem. Fig. 1 illustrates an example of AR and Parallel AR solution construction.
In this work, we propose a model architecture, PARCO, to solve such cooperative multi-agent CO problems efficiently. Drawing on studies on multi-agent RL (MARL) which promotes parameter sharing in cooperative MARL settings with a shared action space (Yu et al., 2022), PARCO establishes a central auto-regressive policy $\pi_{\theta}:\mathcal{S}\rightarrow P(\mathcal{A}_{1}\times...\times\mathcal{A}_{m})$ to decode the next action for the $m$ agents in parallel. This not only makes PARCO agnostic to the number of agents prevalent in the respective CO problem, but also enhances both solution quality and construction speed.
4.2 Parallel Autoregressive Model
In this section, we present our general architecture to solve CO problems efficiently cast as multi-agent sequential decision-making problems. PARCO follows the general encoder-decoder architecture prevalent in autoregressive NCO (Kool et al., 2018; Bello et al., 2016; Vinyals et al., 2015).
Figure 2: Overview of the PARCO model.
Encoder
The job of the encoder in autoregressive NCO approaches is to learn a mapping from the graph representing the CO problem to a higher dimensional embedding space. This typically involves a problem depending on the initial embedding layer, projecting the graph nodes using their feature representations. This is also the case for PARCO, whose encoder first maps features of the graph instance to an embedding space to learn deep relationships between the nodes through a stack of transformer blocks, similar to Kool et al. (2018). The core of each transformer block is a multi-head attention (MHA) layer, which enables message passing between the nodes in the graph and can be defined as follows (Vaswani et al., 2017):
$\text{MHA}(Q,K,V)$ $=\left(\mathop{\big{\|}}_{i=1}^{h}\text{Attn}(QW^{Q}_{i},KW^{K}_{i}VW^{V}_{i})\right)W^{O}$
where
$\text{Attn}(Q,K,V)$ $=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V\,.$
[…]
PARCO additionally defines an agent projection layer that maps agents into the same embedding space as the problem nodes, whose exact definition heavily depends on the problem. In routing problems, for example, agent embeddings can be represented by the depot location they are starting from and their capacities, while an agent in scheduling problems can be represented by the machine whose schedule it is constructing. Lastly, encoded actions and agents can further be enhanced by performing cross-attention between them, depending on the structure of the problem (Kwon et al., 2021). We describe the specific encoder details for the different problems covered in this work in the Appendix.
Having a separate agent encoder contrasts other multi-agent (NCO) approaches like Zhang et al. (2020b), which use separate policies $\pi^{k}_{\theta_{m}}$ for each agent. This has a substantial drawback, as a model trained for a given number of agents cannot be used to solve problems with a different number of agents. The PARCO encoder, on the other hand, can generate representations for arbitrary agents, making it agnostic to the problem size.
Decoder
The decoder is the core of the PARCO architecture. We adopt the pointer mechanism from Kool et al. (2018), which yields unnormalized log-probabilities $\bm{z}\in\mathbb{R}^{N}$ for each action/node:
$\bm{z}=C\cdot tanh\left(\frac{\bm{u}^{\top}L}{\sqrt{d_{h}}}\right)$ (3)
where $C$ is a scale parameter, $L$ is a projection of the node embeddings $\bm{h}$ and $\bm{u}$ is the output of an MHA layer, in which the query is defined by a single context embedding $\bm{q}^{c}_{t}\in\mathbb{R}^{d_{h}}$, specifying the current state of the problem at decoding step $t$ and keys and values are the node embeddings from the encoder. Each attention head of the MHA layer can further fuse the projection of the encoder embeddings with dynamic embeddings representing current information about the particular node, i.e.
$\text{Attn}(\bm{q}^{c}_{t}W^{Q}_{i},\,\bm{h}W^{K}_{i}+W^{K}_{\delta}\delta_{t},\,\bm{h}W^{V}_{i}+W^{V}_{\delta}\delta_{t})$ (4)
with $\delta_{t}$ encapsulating the dynamic elements of the problem nodes and $W^{K}_{\delta},W^{V}_{\delta}\in\mathbb{R}^{|\delta|\times d_{m}}$ projecting them into embedding space.
For PARCO we introduce a novel Multiple Pointer Mechanism based on Eq. 3 of the following form:
$Z=C\cdot tanh\left(\frac{UL^{\top}}{\sqrt{d_{h}}}\right)$ (5)
Here, $U\in\mathbb{R}^{m\times d_{h}}$ is the output of an equivalent MHA layer than the one used in Kool et al. (2018). However, instead of a single context embedding, PARCO uses the $m$ agent embeddings of the encoder, enriched by the current state’s context and individual agent’s state embeddings, as queries $Q^{c}_{t}\in\mathbb{R}^{m\times d_{h}}$. The resulting $Z\in\mathbb{R}^{m\times N}$ log-probabilities are used to sample $m$ actions - one for each agent - in parallel.
Communication Layers
To enable effective agent coordination, we introduce a communication layer before the Multiple Pointer Mechanism, which updates the agent embeddings through messages exchanged with other agents. This component is vital for parallel decoding of multiple actions, as the quality of an action chosen by one agent is significantly influenced by how it affects the actions of other agents.
Given the $m$ agent queries at decoding step $t$, the communication layer applies a transformer block (Vaswani et al., 2017) to capture intra-agent relationships and spot potential conflicts between agents:
$H^{\prime}$ $=\text{Norm}(\text{MHA}(Q^{c}_{t},Q^{c}_{t},Q^{c}_{t})+Q^{c}_{t})$ (6)
$Q^{\prime}_{t}$ $=\text{Norm}(\text{MLP}(H^{\prime})+H^{\prime})$ (7)
where Norm denotes a normalization layer (Ioffe and Szegedy, 2015; Zhang and Sennrich, 2019) and MLP represents the multi-layer perceptron. Self-attention in the MHA layer enables message passing between agents based on their embeddings. By applying the communication layer after fusing the agent embeddings with the dynamic context embeddings, it can capture dynamic relationships that could not be resolved during encoding, thereby enabling the decoder to make more informed decisions. The updated query $Q^{\prime}_{s}$ is finally passed into the Multiple Pointer Mechanism.
4.2.1 Conflicts Handlers
When sampling from the probability distribution generated by the Multiple Pointer Mechanism, it is possible for multiple agents to select the same node simultaneously, resulting in a conflict. These conflicts must be resolved to ensure the generation of a feasible solution. To address this, we introduce a Priority-based Conflict Handler $\mathcal{H}$, which resolves such conflicts by leveraging a predefined priority scheme. The conflict handler can be defined as a function $\mathcal{H}:\mathbb{N}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{N}^{m}$ with number of agents $m$. The priorities are determined based on the current actions and states. Given an input vector of actions $\bm{a}=(a_{1},a_{2},\dots,a_{m})$ where $a_{i}\in\mathbb{N}$, the corresponding priority vector $\bm{p}=(p_{1},p_{2},\dots,p_{m})$ where $p_{i}\in\mathbb{R}$ and the fallback actions (i.e., the previous node) $\bm{r}=(r_{1},r_{2},\dots,r_{m})$ where $r_{i}\in\mathbb{N}$, the output is another vector of actions $\bm{a}^{\prime}=(a^{\prime}_{1},a^{\prime}_{2},\dots,a^{\prime}_{m})$ such that any conflicts are resolved based on the priority order as illustrated in Algorithm 1.
Algorithm 1 Priority-based Conflict Handler
0:  Actions $\bm{a}\in\mathbb{N}^{m}$, Priorities $\bm{p}\in\mathbb{R}^{m}$, Fallback actions $\bm{r}\in\mathbb{N}^{m}$
0:  Resolved Actions $\bm{a^{\prime}}\in\mathbb{N}^{m}$
1:  $\sigma\leftarrow\text{argsort}(\bm{p},\text{descending}=\text{True})$Sort indices by priority}
2:  $\hat{\bm{a}}\leftarrow\bm{a}[\sigma]$Reorder actions according to priority}
3:  Initialize conflict mask $M\leftarrow\bm{0}^{m}$
4:  for $i=2$ to $m$ do
5:     if $\hat{\bm{a}}_{i}\in\{\hat{\bm{a}}_{1},\ldots,\hat{\bm{a}}_{i-1}\}$ then
6:        $M_{i}\leftarrow 1$Identify conflicts}
7:     end if
8:  end for
9:  $\hat{\bm{a}}\leftarrow(1-M)\odot\hat{\bm{a}}+M\cdot r$Resolve conflicts by assigning fallback action $r$}
10:  $\bm{a}^{\prime}\leftarrow\hat{\bm{a}}[\sigma^{-1}]$Reorder actions back to original sequence}
11:  return  $\bm{a^{\prime}}$
4.2.2 Step Definition
We define a single parallel step of the model similarly to Eq. 1 as follows:
$p_{\theta}(\bm{a}|\bm{x})$ $=\prod_{i=1}^{m}g_{\theta}(a_{t,i}^{\prime}|a_{t-1,i}^{\prime},\dots,a_{0,i}^{\prime},\bm{h})$ (8)
Here, $a_{t,i}^{\prime}$ denotes the action executed by agent $i$ at time $t$, after passing through the conflict handler $\mathcal{H}$:
$a_{t,i}^{\prime}=\mathcal{H}(a_{t,1},a_{t,2},\dots,a_{t,m};\bm{p}_{t},\bm{r}_{t}),$
where $\bm{p}_{t}$ and $\bm{r}_{t}$ are priorities and fallback actions, respectively, at the current decoding step $t$. The conflict handler $\mathcal{H}$ ensures that multiple agents do not select the same node, thus maintaining the integrity and feasibility of the solution.
4.3 Training
PARCO is a centralized multi-agent sequential decision-making model, with a shared policy $\pi_{\theta}$ for all agents and a global reward $R$. Thus, PARCO can be trained using any of the training algorithms adopted in the single-agent NCO literature. We train PARCO via the REINFORCE gradient estimator (Williams, 1992) with a shared baseline as outlined by Kwon et al. (2020) and Kim et al. (2022):
$\nabla_{\theta}\mathcal{L}\approx\frac{1}{B\cdot L}\sum_{i=1}^{B}\sum_{j=1}^{L}G_{ij}\nabla_{\theta}\,\text{log}\,p_{\theta}(A_{ij}|\bm{x}_{i})$ (9)
where $B$ is the size of the mini-batch and $G_{ij}$ is the advantage $R(A_{ij},\bm{x}_{i})-b^{\text{shared}}(\bm{x}_{i})$ of a solution $A_{ij}$ compared to the shared baseline $b^{\text{shared}}_{i}$ of problem instance $\bm{x}_{i}$.
5 Experiments
In this section, we present the experimental results of PARCO in two routing problems, the min-max heterogenous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP), and a scheduling problem, namely the flexible flow shop problem (FFSP). We provide more details about the problem and experimental setups in Appendix A and Appendix B respectively.
5.1 Problem Descriptions
HCVRP
The min-max HCVRP consists of $m$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot. The goal is to minimize the makespan, i.e., the worst route. Baselines include SISR (Christiaens and Vanden Berghe, 2020), Genetic Algorithm (GA) (Karakatič and Podgorelec, 2015), Simulated Annealing (SA) (İlhan, 2021), the Attention Model (AM) (Kool et al., 2018), Equity Transformer (ET) (Son et al., 2024), the model from Li et al. (2022) (DRLLi), and the state-of-the-art neural baseline 2D-Ptr (Liu et al., 2024c).
OMDCPDP
The OMDCPDP problem is a practical variant of the pickup and delivery problem in which agents have a stacking limit of orders that can be carried at any given time. Pickup and delivery locations are paired, and pickups must be visited before deliveries. Multiple agents start from different depots with no need to go back (open). The goal is to minimize the sum of arrival times to delivery locations, i.e. minimizing the cumulative lateness. We include ORTools (Furnon and Perron, 2024) as a classical baseline, the Heterogeneous Attention Model (HAM) (Li et al., 2021a) for sequential decision-making and MAPDP (Zong et al., 2022) for parallel decision-making.
FFSP
In the flexible flow shop problem (FFSP), $N$ jobs must be processed across $S$ stages, each with multiple machines ($m>1$). Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously. The goal is to schedule the jobs so that all jobs are finished in the shortest time possible. Notable benchmarks include the MatNet model (Kwon et al., 2021), the Random and Shortest Job First (SJF) dispatching rules, as well as the evolutionary algorithms Particle Swarm Optimization (PSO), and Genetic Algorithm (GA) (Hejazi and Saghafian, 2005).
Table 1: Benchmarks and results of our model for min-max HCVRP problems of varying sizes and agent numbers. Highlighting cost ($\downarrow$) and average gaps ($\downarrow$) from the best-known solutions of classical heuristic solvers. Inference times are shown in seconds in parentheses ()($\cdot$)( ⋅ ). (g.) refers to greedy performance while (s.) refers to sampling 1280 solutions.
[…]
$m$ 3 5 7 3 5 7 3 5 7 avg.
SISR 6.57 (271) 4.00 (274) 2.91 (276) 8.52 (425) 5.10(430) 3.69 (434) 10.29 (615) 6.17 (623) 4.45 (625) 0.00
GA 9.21 (233) 6.89 (320) 5.98 (405) 12.32 (347) 8.95 (465) 7.58 (578) 15.33 (479) 10.93 (623) 9.10 (772) 74.90
SA 7.04 (130) 4.39 (289) 3.30 (362) 9.17 (318) 5.61 (417) 4.17 (515) 11.13 (434) 6.80 (557) 5.01 (678) 10.21
AM (g.) 8.49 (0.08) 5.51 (0.08) 4.15 (0.09) 10.81 (0.10) 6.87 (0.11) 5.18 (0.10) 12.68 (0.14) 8.10 (0.13) 6.13 (0.13) 33.80
ET (g.) 7.58 (0.15) 4.76 (0.17) 3.58 (0.16) 9.76 (0.21) 6.01 (0.20) 4.43 (0.23) 11.74 (0.25) 7.25 (0.25) 5.23 (0.26) 17.67
DRLLi (g.) 7.43 (0.19) 4.71 (0.22) 3.60 (0.25) 9.64 (0.25) 5.97 (0.30) 4.52 (0.33) 11.44 (0.32) 7.06 (0.37) 5.38 (0.43) 17.08
2D-Ptr (g.) 7.20 (0.11) 4.48 (0.11) 3.31 (0.11) 9.24 (0.15) 5.65 (0.15) 4.14 (0.14) 11.12 (0.18) 6.75 (0.18) 4.92 (0.17) 10.54
PARCO (g.) 7.12 (0.12) 4.40 (0.11) 3.25 (0.11) 9.14 (0.15) 5.53 (0.15) 4.04 (0.15) 10.98 (0.19) 6.61 (0.18) 4.79 (0.18) 8.56
AM (s.) 7.62 (0.14) 4.82 (0.13) 3.63 (0.14) 9.92 (0.20) 6.19 (0.21) 4.64 (0.22) 11.82 (0.29) 7.45 (0.28) 5.58 (0.28) 20.69
ET (s.) 7.14 (0.21) 4.46 (0.22) 3.33 (0.22) 9.19 (0.30) 5.64 (0.30) 4.17 (0.31) 11.20 (0.41) 6.85 (0.38) 4.98 (0.40) 10.89
DRLLi (s.) 6.97 (0.30) 4.34 (0.36) 3.25 (0.43) 9.10 (0.45) 5.54 (0.55) 4.13 (0.65) 10.90 (0.60) 6.65 (0.76) 4.98 (0.92) 8.78
2D-Ptr (s.) 6.82 (0.13) 4.20 (0.13) 3.09 (0.14) 8.85 (0.17) 5.36 (0.18) 3.90 (0.19) 10.71 (0.22) 6.46 (0.23) 4.68 (0.24) 4.84
PARCO (s.) 6.82 (0.31) 4.17 (0.31) 3.06 (0.31) 8.79 (0.36) 5.28 (0.36) 3.83 (0.35) 10.61 (0.40) 6.36 (0.40) 4.58 (0.40) 3.65
5.2 Experimental Setup
We perform all experiments on a machine equipped with two Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz CPUs with a total $128$ threads and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM. Training runs of PARCO take less than 24 hours each. During inference, we employ only one CPU and a single GPU. We report key metrics such as solution cost, inference times, and gaps in best-known solutions. We keep most settings of PARCO consistent across experiments, i.e., we use a Communication Layer and Priority-based Conflict Handling based on the highest probability action from the model output. We provide additional details regarding training and testing setups in the Appendix.
(a) Effect of Communication Layers.
(b) Effect of Conflict Handlers.
(c) PARCO vs AR decoding speed.
Figure 3: Ablation studies on PARCO components. PARCO scales better than AR models as ET.
Table 2: Benchmarks and results of our model for OMDCPDP problems of varying sizes and agent numbers.
[…]
$N$ 50 100 200 1000 avg.
$m$ 10 18 25 20 35 50 40 70 100 100 250 500
OR-Tools 23.73 18.64 16.92 43.34 34.80 31.89 78.35 64.93 60.16 562.11 333.01 310.55 0.00%
HAM (g.) 34.85 24.79 18.52 65.95 51.56 35.41 124.16 98.32 67.45 740.34 475.76 326.07 33.52%
MAPDP (g.) 27.95 20.13 17.54 52.44 38.23 $3$ 10.80%
PARCO (g.) 24.70 19.19 17.44 44.88 35.51 32.79 81.14 66.95 61.95 531.95 344.67 302.09 1.96%
HAM (s.) 32.84 25.45 18.12 63.59 49.12 34.90 122.77 96.32 67.01 739.77 471.09 321.50 31.02%
MAPDP (s.) 25.49 20.02 17.17 49.55 37.37 33.25 - - - - - - 7.04%
PARCO (s.) 24.27 18.71 17.00 43.36 34.82 32.19 79.23 65.83 61.12 526.52 338.78 298.71 -0.01%
Table 3: FFSP results. PARCO improves the MatNet performance and is also more than 4×4$\times$4 × faster overall.
Method FFSP20 FFSP50 FFSP100
MS Gap Time (training) MS Gap Time (training) MS Gap Time (training)
CPLEX 46.4 21.0 60s - -
CPLEX 36.6 6.4 600s - -
Random 47.8 22.9 0.1s 93.2 44 0.2s 167.2 78.0 0.3s
Shortest Job First 31.3 6.4 0.1s 57.0 7.8 0.1s 99.3 10.1 0.2s
Genetic Algorithm 30.6 5.7 25.4s 56.4 7.2 57.8s 98.7 9.5 104.5s
Particle Swarm Opt. 29.1 4.2 46.8s 55.1 5.9 92.9s 97.3 8.1 173.4s
MatNet 27.26 3.0 0.9s (2h) 51.52 0.5 2.1s (5.25h) 91.58 0.0 4.9s (28h)
PARCO (w/o CL) 27.02 2.1 0.2s (0.25h) 52.29 2.0 0.4s (1.25h) 92.15 2.0 0.9s (5.25h)
PARCO 26.47 0.0 0.2s (0.5h) 51.25 0.0 0.5s (2h) 91.55 0.0 1.1s (6.25h)
5.3 Experimental Results
Main experiments
The main experiments showcasing the performance of PARCO are in Table 1, Table 2, Table 3 for HCVRP, OMDCPDP, and FFSP, respectively. Our PARCO consistently outperforms SotA neural baselines across a variety of experiments. We additionally note another property of PARCO: unlike neural baselines in HCVRP and MAPDP in the OMDCPDP, which are trained specifically for a single size and number of agents, PARCO is a single model trained on multiple location and agent distributions at the same time; nonetheless, our single model can outperform baselines trained ad hoc on specific distributions, demonstrating our method’s flexibility.
Generalization
We additionally study the generalization performance in Table 2 on the right, which are the sizes and number of agents unseen during training in OMDCPDP. Unlike MAPDP, which is constrained to a specific number of agents, PARCO can successfully generalize to unseen sizes and $m$. Notably, PARCO can even outperform Google ORTools for larger-scale instances, demonstrating remarkable scalability.
Table 4: Effect of different ##\## of agents in FFSP.
$J\times S\times m$ Metric MatNet PARCO
$50\times 3\times 6$ Obj. 35.69 33.58
Gap 6.3% 0.0%
Duration 1.1h 10.3m
Avg. Steps 200.05 19.10
$50\times 3\times 8$ Obj. 28.59 25.62
Gap 11.6% 0.0%
Duration 1.5h 9.4m
Avg. Steps 205.95 16.80
$50\times 3\times 10$ Obj. 25.05 21.46
Gap 16.7% 0.0%
Duration 1.8h 7.5m
Avg. Steps 215.85 13.45
Effect of Communication Layers
We showcase the importance of Communication Layers in Fig. 3 (a). We benchmark 1) No Communication (only context features), 2) communication via an MLP, 3) communication via an MHA layer, and 4) Our Communication layer. Our Communication Layer consistently outperforms other methods.
Effect of Conflict Handlers
Fig. 3 (b) shows the effect of the Priority-based Conflict Handler with different priorities methods $\bm{p}$ in the OMDCPDP. We consider the following: 1) Random: the priority is chosen randomly as in Zong et al. (2022), 2) First: the priority is chosen by the agent index $k=1,\dots,m$, 3) Smallest: gives priority to the agent that has traveled the least, 4) Closest: priority is given to the closest agent and 5) High Probability: priority is given to the agent whose probability of selecting the conflicting action is the highest. High Probability can consistently outperform other methods. We note that this is also the most general and problem-agnostic since methods 2-4 are problem-specific.
Scalability
Finally, we showcase PARCO’s scalability in terms of speed in large-scale instances in Fig. 3 (c) compared to the constructive autoregressive ET (Son et al., 2024). Interestingly, PARCO’s gap grows with more agents. This is because our method can fully exploit agent parallelism and can thus be a strong candidate for large-scale real-time CO applications.
6 Conclusion
In this paper, we introduced a novel approach, PARCO - Parallel AutoRegressive Combinatorial Optimization - that utilizes parallel autoregressive (AR) policies to tackle multi-agent NCO problems. We introduced a new Communication Layer to allow for the agents to effectively coordinate their next steps during decoding, alongside a Multiple Pointer Mechanism coupled with a Priority-based Conflicts Handler that can generate feasible solutions efficiently. We validate the effectiveness of PARCO through extensive experiments with different CO problems from the routing and scheduling domain and demonstrate its competitive performance against classical heuristic and neural baselines.
While PARCO is already faster than most evaluated neural baselines and decreases the number of required decoding steps by more than 90% in FFSP50 instances with 10 agents as shown in Table 4, it still does not achieve the potential reduction from $\mathcal{O}(N)$ to $\mathcal{O}(\frac{N}{m})$ in many cases. This is mainly attributed to the fact that agents can have conflicts that we currently resolve with the priority-based conflict handler. Giving one agent precedence and forcing others to stay in their current position results in more than necessary decoding steps. In future work, we will try to address this limitation by constructing a custom loss function, which penalizes conflicts and makes agents learn to avoid them during training. Further, we plan to extend our work with learnable improvement methods to enable fast first solutions coupled with powerful local search. Given PARCO’s solution quality and efficiency, a further exciting avenue of future research is to couple our model with population-based approaches (Grinsztajn et al., 2022; Chalumeau et al., 2023; Hottung et al., 2024) that could search the latent space for diverse and better solutions.
Acknowledgements
We gratefully acknowledge OMELET for supporting us with computing. We also thank people in the AI4CO open research community who have contributed to RL4CO, which this work is based on.
Appendix
[…]
Appendix A Problem Descriptions
A.1 HCVRP
A.1.1 Problem Definition
The min-max HCVRP (Heterogeneous Capacitated Vehicle Routing Problem) consists of $m$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot. The goal is to minimize the makespan, i.e., the worst route.
A.1.2 Mathematical Formulation
Consider a problem with $N+1$ nodes (including $N$ customers and a depot) and $m$ vehicles. The depot is indexed as 0, and customers are indexed from 1 to $N$.
Indices
$i,j$ Node indices, where $i,j=0,\ldots,N$ (0 represents the depot)
$k$ Vehicle index, where $k=1,\ldots,m$
Parameters
$N$ Number of customer nodes (excluding depot)
$m$ Number of vehicles
$X_{i}$ Location of node $i$
$d_{i}$ Demand of node $i$ ($d_{0}=0$ for the depot)
$Q_{k}$ Capacity of vehicle $k$
$f_{k}$ Speed of vehicle $k$
$c_{ij}$ Distance between nodes $i$ and $j$
Decision Variables
$x_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ travels directly from node }i\text{ to node }j\\ 0&\text{otherwise}\end{array}\right.$
$l_{ijk}$ Remaining load of vehicle $k$ before travelling from node $i$ to node $j$
Objective Function:
$\min\max_{k=1,\ldots,m}\left(\sum_{i=0}^{N}\sum_{j=0}^{N}\frac{c_{ij}}{f_{k}}x_{ijk}\right)$ (10)
Subject to:
$\sum_{k=1}^{m}\sum_{j=0}^{N}x_{ijk}$ $=1$ $i=1,\ldots,N$ (11)
$\sum_{i=0}^{N}x_{ijk}-\sum_{h=0}^{N}x_{jhk}$ $=0$ $j=0,\ldots,N,\;k=1,\ldots,m$ (12)
$\sum_{k=1}^{m}\sum_{i=0}^{N}l_{ijk}-\sum_{k=1}^{m}\sum_{h=0}^{N}l_{jhk}$ $=d_{j}$ $j=1,\ldots,N$ (13)
$d_{j}x_{ijk}\leq l_{ijk}$ $\leq(Q_{k}-d_{i})\cdot x_{ijk}$ $i,j=0,\ldots,N,\;k=1,\ldots,m$ (14)
$x_{ijk}$$\in\{0,1\}$ $i,j=0,\ldots,N,\;k=1,\ldots,m$ (15)
$l_{ijk}\geq 0,d_{i}$ $\geq 0$ $i,j=0,\ldots,N,\;k=1,\ldots,m$ (16)
Constraint Explanations: The formulation is subject to several constraints that define the feasible solution space. Equation (11) ensures that each customer is visited exactly once by one vehicle. The flow conservation constraint (12) guarantees that each vehicle that enters a node also leaves that node, maintaining route continuity. Demand satisfaction is enforced by constraint (13), which ensures that the difference in load before and after serving a customer equals the customer’s demand. The vehicle capacity constraint (14) ensures that the load carried by a vehicle does not exceed its capacity and is sufficient to meet the next customer’s demand.
A.2 OMDCPDP
A.2.1 Problem Definition
The OMDCPDP (Open Multi-Depot Capacitated Pickup and Delivery Problem) is a practical variant of the pickup and delivery problem in which agents have a stacking limit of orders that can be carried at any given time. Pickup and delivery locations are paired, and pickups must be visited before deliveries. Multiple agents start from different depots without returning (open). The goal is to minimize the sum of arrival times to delivery locations, i.e., minimizing the cumulative lateness.
A.2.2 Mathematical Formulation
Indices
$i,j$ Node indices, where $i,j=1,\ldots,2N$
$k$ Vehicle index, where $k=1,\ldots,m$
Sets
$P$ Set of pickup nodes, $P=\{1,\ldots,N\}$
$D$ Set of delivery nodes, $D=\{N+1,\ldots,2N\}$
Parameters
$N$ Number of pickup-delivery pairs
$m$ Number of vehicles
$c_{ij}$ Travel time between nodes $i$ and $j$
$Q_{k}$ Capacity (stacking limit) of vehicle $k$
$o_{k}$ Initial location (depot) of vehicle $k$
Decision Variables
$x_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ travels directly from node }i\text{ to node }j\\ 0&\text{otherwise}\end{array}\right.$
$y_{ik}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ visits node }i\\ 0&\text{otherwise}\end{array}\right.$
$t_{i}$ Arrival time at node $i$
$l_{ik}$ Load of vehicle $k$ after visiting node $i$
Objective Function:
$\min\sum_{i=N+1}^{2N}t_{i}$ (17)
Subject to:
$\sum_{k=1}^{m}y_{ik}$ $=1$ $i=1,\ldots,2N$ (18)
$\sum_{j=1}^{2N}x_{o_{k},j,k}$ $=1$ $k=1,\ldots,m$ (19)
$\sum_{i=1}^{2N}x_{ijk}-\sum_{h=1}^{2N}x_{jhk}$ $=0$ $j=1,\ldots,2N,\;k=1,\ldots,m$ (20)
$y_{ik}$ $=\sum_{j=1}^{2N}x_{ijk}$ $i=1,\ldots,2N,\;k=1,\ldots,m$ (21)
$t_{i}+c_{ij}-M(1-x_{ijk})$ $\leq t_{j}$ $i,j=1,\ldots,2N,\;k=1,\ldots,m$ (22)
$t_{i}$ $\leq t_{i+N}$ $i\in P$ (23)
$l_{ik}+1-M(1-x_{ijk})$ $\leq l_{jk}$ $i\in P,j\neq i+N,k=1,\ldots,m$ (24)
$l_{ik}-1+M(1-x_{ijk})$ $\geq l_{jk}$ $i\in D,j\neq i-N,k=1,\ldots,m$ (25)
$0\leq l_{ik}$ $\leq Q_{k}$ $i=1,\ldots,2N,\;k=1,\ldots,m$ (26)
$x_{ijk},y_{ik}$$\in\{0,1\}$ $i,j=1,\ldots,2N,\;k=1,\ldots,m$ (27)
$t_{i}$ $\geq 0$ $i=1,\ldots,2N$ (28)
Constraints Explanations: Equation (18) ensures that each node is visited exactly once. Constraint (19) guarantees that each vehicle starts from its designated depot. The flow conservation constraint (20) ensures route continuity for each vehicle. Equation (21) defines the relationship between x and y variables. Time consistency is enforced by constraint (22), while (23) ensures that pickups are visited before their corresponding deliveries. Constraints (24) and (25) manage the load changes during pickup and delivery operations. Finally, the vehicle capacity constraint (26) ensures that the load never exceeds the vehicle’s stacking limit.
A.3 FFSP
A.3.1 Problem Definition
The flexible flow shop problem (FFSP) is a challenging and extensively studied optimization problem in production scheduling, involving $N$ jobs that must be processed across $i=1\dots S$ stages, each with multiple machines ($m_{i}>1$). Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously. The FFSP can naturally be viewed as a multi-agent CO problem by considering each machine as an agent that constructs its own schedule. Adhering to autoregressive CO, agents construct the schedule sequentially, selecting one job (or no job) at a time. The job selected by a machine (agent) at a specific stage in the decoding process is scheduled at the earliest possible time, that is, the maximum of the time the job becomes available in the respective stage (i.e., the time the job finished on prior stages) and the machine becoming idle. The process repeats until all jobs for each stage have been scheduled, and the ultimate goal is to minimize the makespan, i.e., the total time required to complete all jobs.
A.3.2 Mathematical Formulation
We use the mathematical model outlined in Kwon et al. (2021) to define the FFSP:
Indices
$i$ Stage index
$j,l$ Job index
$k$ Machine index in each stage
Parameters
$N$ Number of jobs
$S$ Number of stages
$m_{i}$ Number of machines in stage $i$
$M$ A very large number
$p_{ijk}$ Processing time of job $j$ in stage $i$ on machine $k$
Decision variables
$C_{ij}$ Completion time of job $j$ in stage $i$
$X_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if job }j\text{ is assigned to machine }k\text{ in stage }i\\ 0&\text{otherwise}\end{array}\right.$
$Y_{ilj}$$\left\{\begin{array}[]{ll}1&\text{if job }l\text{ is processed earlier than job }j\text{ in stage }i\\ 0&\text{otherwise}\end{array}\right.$
Objective:
$\min\left(\max_{j=1..n}\{C_{Sj}\}\right)$ (29)
Subject to:
$\sum_{k=1}^{m_{i}}X_{ijk}$ $=1$ $i=1,\dots,S;\;j=1,\dots,N$ (30)
$Y_{iij}$ $=0$ $i=1,\dots,S;\;j=1,\dots,N$ (31)
$\sum_{j=1}^{N}\sum_{l=1}^{N}Y_{ilj}$ $=\sum_{k=1}^{m_{i}}\max\left(\sum_{j=1}^{n}(X_{ijk}-1),0\right)$ $i=1,\dots,S$ (32)
$Y_{ilj}$ $\leq\max\left(\max_{k=1\dots m_{i}}\{X_{ijk}+X_{ilk}\}-1,0\right)$ $i=1,\dots,S;\;j,l=1,2,\dots,N$ (33)
$\sum_{l=1}^{N}Y_{ilj}$ $\leq 1$ $i=1,2,\dots,S;\;j=1,2,\dots,N$ (34)
$\sum_{j=1}^{N}Y_{ilj}$ $\leq 1$ $i=1,2,\dots,S;\;l=1,2,\dots,N$ (35)
$C_{1j}$ $\geq\sum_{k=1}^{m_{1}}p_{1jk}\cdot X_{1jk}$ $j=1,2,\dots,N$ (36)
$C_{ij}$ $\geq C_{i-1j}+\sum_{k=1}^{m_{i}}p_{ijk}\cdot X_{ijk}$ $i=2,3,\dots,S;\;j=1,2,\dots,N$ (37)
$C_{ij}+M(1-Y_{ilj})$ $\geq C_{il}+\sum_{k=1}^{m_{i}}p_{ijk}\cdot X_{ijk}$ $i=1,2,\dots,S;\;j,l=1,2,\dots,N$ (38)
Constraint Explanations: Here, the objective function Eq. 29 minimizes the makespan of the resulting schedule, that is, the completetion time of the job that finishes last. The schedule has to adhere to several constraints: First, constraint set Eq. 30 ensures that each job is assigned to exactly one machine at each stage. Constraint sets Eq. 31 through Eq. 35 define the precedence relationships between jobs within a stage. Specifically, constraint set Eq. 31 ensures that a job has no precedence relationship with itself. Constraint set Eq. 32 ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned. Constraint set Eq. 33 dictates that precedence relationships can only exist among jobs assigned to the same machine. Additionally, constraint sets Eq. 34 and Eq. 35 restrict a job to having at most one preceding job and one following job.
Moving on, constraint set Eq. 36 specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage. The relationship between the completion times of a job in consecutive stages is described by constraint set Eq. 37. Finally, constraint set Eq. 38 ensures that no more than one job can be processed on the same machine simultaneously.
Appendix B Experimental Details
B.1 HCVRP
B.1.1 Baselines
SISR
The Slack Induction by String Removals (SISR) approach (Christiaens and Vanden Berghe, 2020) offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process. It combines techniques for route dismantling and reconstruction, along with vehicle fleet minimization strategies. SISR is applied across various VRP scenarios, including those with specific pickup and delivery tasks. In our experiments, we adhere to the hyperparameters provided in the original paper with $\bar{c}=10,L^{\mathrm{max}}=10,\alpha=10^{-3},\beta=10^{-2},T_{0}=100,T_{f}=1,\mathrm{iter}=3\times 10^{5}\times N$.
GA
The Genetic Algorithm (GA) (Karakatič and Podgorelec, 2015) is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes. Particularly effective in Multi-Depot Vehicle Routing Problems (MDVRP), GA quickly generates adequate solutions with reasonable computational resources. In our experiment, we follow the same carefully tuned hyperparameters from Liu et al. (2024c) with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
SA
The Simulated Annealing (SA) method (İlhan, 2021) targets the capacitated vehicle routing problem (CVRP) using a population-based approach combined with crossover operators. It incorporates local search and the improved 2-opt algorithm to refine routes alongside crossover techniques to speed up convergence. In our experiment, we follow the same carefully tuned hyperparameters from Liu et al. (2024c) with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
AM
The Attention Model (AM) (Kool et al., 2018) applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems. It utilizes attention layers for model improvement and trains using REINFORCE with deterministic rollouts. In our studies, we adopt adjustments from the DRLLi framework, which involves selecting vehicles sequentially and then choosing the next node for each. Additionally, vehicle-specific features are incorporated into the context vector generation to distinguish between different vehicles.
ET
The Equity-Transformer (ET) approach (Son et al., 2024) addresses large-scale min-max routing problems by employing a sequential planning approach with sequence generators like the Transformer. It focuses on equitable workload distribution among multiple agents, applying this strategy to challenges like the min-max multi-agent traveling salesman and pickup and delivery problems. In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP, and integrate vehicle features into both the input layer and the context encoder.
DRLLi
The DRL approach for solving HCVRP by Li et al. (2022) employs a transformer architecture similar to Kool et al. (2018) in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions. We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents $m$, each with always the same characteristics.
2D-Ptr
The 2D Array Pointer network (2D-Ptr) (Liu et al., 2024c) addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively. This approach facilitates dynamic, real-time decision-making for route optimization. Its decoder employs a 2D array pointer for action selection, prioritizing actions over vehicles. The model is designed to adapt to changes in vehicle and customer numbers, ensuring robust performance across different scenarios. For our study, 2D-Ptr was configured following the original study’s parameters for consistent comparison with 8 heads, 128 hidden dimensions for model structure; $\mathrm{lr}=10^{-4},\lambda_{\mathrm{Adam}}=0.995,\kappa_{\mathrm{L2}}=3$ for turning; 50 epochs, 2500 batches per epoch, 512 instances per batch for training.
B.1.2 Datasets
Train data generation
Neural baselines were trained with the specific number of nodes $N$ and number of agents $m$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(60,100)$ and $m\sim\mathcal{U}(3,7)$. As we show in the Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. The depot location is similarly sampled using the same uniform distribution. The demand $d_{i}$ for each customer $i$ is also drawn from a uniform distribution $\mathcal{U}(1,10)$, with the depot having no demand, i.e., $d_{0}=0$. Each vehicle $k$, where $k=1,\ldots,m$, is assigned a capacity $Q_{k}$ sampled from a uniform distribution $\mathcal{U}(20,41)$. The speed $f_{k}$ of each vehicle is uniformly distributed within the range $\mathcal{U}(0.5,1.0)$.
Testing
Testing is performed on the 1280 instances per test setting from Liu et al. (2024c). In Table 1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward). We report optimality gaps to ORTools and solution time in seconds in parentheses.
B.1.3 PARCO Network Hyperparameters
Encoder
Initial Embedding. This layer projects initial raw features to hidden space. For the depot, the initial embedding is the positional encoding of the depot’s location $X_{0}$. For agents, the initial embedding is the encoding for the initial location, capacity, and speed. Main Encoder. we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
Decoder
Context Embedding. This layer projects dynamic raw features to hidden space. The context is the embedding for the depot states, current node states, current time, remaining capacities, time of backing to the depot, and number of visited nodes. These features are then employed to update multiple queries $q_{k},~{}k=1,\dots,m$ simultaneously. Main Decoder. Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
Communication Layer
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP. Unlike the encoder layer, which acts between all $m+N$ nodes, communication layers are lighter because they communicate between $m$ agents.
Agent Handler
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action $a$ is the highest. In other words, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
B.1.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma and Ba, 2014) with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $4\times 10^{5}$ randomly generated data. Training takes around 15 hours in our configuration.
B.2 OMDCPDP
B.2.1 Datasets
Train data generation
Neural baselines were trained with the specific number of nodes $2N$ and number of agents $m$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $2N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(10,50)$. As we show in the Table 2, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,2N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. Similarly, we sample $m$ initial vehicle locations from the same distribution. We set the demand $d_{i}$ for each customer to $1$ and the capacity of each vehicle to $3$. This emulates realistic settings in which a single package per customer will be picked up and delivered.
Testing
Testing is performed on 100 new instances for each setting of $N$ and $m$ in Table 2 with the distributions from the training settings. We finally test PARCO on large-scale, real-world data from complex, realistic distributions with $32$ instances for each vehicle, pickup, and delivery location in Seoul City, South Korea. This dataset is provided in our source code. Fig. 4 provides an example of such data. In Table 2, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
Figure 4: Real-world instance for the OMDCPDP problem in Seoul City, South Korea, with $N=1000$ locations and $m=100$ agents (green) showing relations (gray) of pickups (red triangles) and their respective deliveries (black crosses).
B.2.2 Baselines
OR-Tools
The OR-Tools (Furnon and Perron, 2024) is open-source software designed to address various combinatorial optimization problems. This toolkit offers a comprehensive suite of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges. Specifically for routing problems like the OMDCPDP, OR-Tools can integrate additional constraints to enhance solution accuracy. Documentation for the Python OR-Tools library, along with example codes for diverse optimization challenges, is readily accessible on their official website. For our experiments, we maintained consistent parameters across various problem sizes and numbers of agents. We configured the global span cost coefficient to $10,000$, selected PATH_CHEAPEST_ARC as the initial solution strategy, followed by GUIDED_LOCAL_SEARCH for local optimization. The solving time was capped at 300 seconds to standardize comparison and measure efficiency.
HAM
The Heterogeneous Attention Model (HAM) (Li et al., 2021a) utilizes a neural network integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes. This approach helps the deep reinforcement learning model to make informed node selections during route planning. We used the same hyperparameter settings from the original HAM experiments.
MAPDP
The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) (Zong et al., 2022) pioneers the use of multi-agent reinforcement learning (MARL) for the cooperative PDP with multiple vehicle agents. This innovative framework introduces a centralized MARL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes. We adapted the MAPDP to fit our OMDCPDP task, utilizing the same encoder as PARCO to ensure a fair comparison. For the decoder and training phases, we kept the same random conflict handler and retained the same hyperparameters as detailed in the original study with 8 heads, 128 hidden dimensions for model structure; $\mathrm{lr}=10^{-3},L_{\mathrm{Adam}}=3$ for turning.
B.2.3 PARCO Network Hyperparameters
Most hyperparameters are kept similar to Section B.1.3; we report all of them nonetheless below.
Encoder
Initial Embedding. This layer projects initial raw features to hidden space. For depots, the initial embeddings encode the location $o_{k}$ and the respective vehicle’s capacity $Q_{k}$. For pickup nodes, the initial embeddings encode the location and paired delivery nodes’ location. For delivery nodes, the initial embeddings encode the location and paired pickup nodes’ location. Main Encoder. We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
Decoder
Context Embedding. This layer projects dynamic raw features to hidden space. The context is the embedding for the depot states $o_{k}$, current node states, current length, remaining capacity, and number of visited nodes. These features are then employed to update multiple queries $q_{k},~{}k=1,\dots,m$ simultaneously. Main Decoder. Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
Communication Layer
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP. Note that unlike the encoder layer, which acts between all $m+N$ nodes, communication layers are lighter because they communicate between $m$ agents.
Agent Handler
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action $a$ is the highest. In other words, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
B.2.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma and Ba, 2014) with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $10^{5}$ randomly generated data. Training takes less than 4 hours in our configuration.
B.3 FFSP
B.3.1 Baselines
MatNet
The comparison of PARCO with other FFSP baselines is shown in table Table 3. Being a neural solver, we benchmark PARCO mainly against MatNet (Kwon et al., 2021), the SotA NCO architecture for the FFSP. MatNet is an encoder-decoder architecture, which is inspired by the attention model (Kool et al., 2018). It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable to encode nodes of different types in a bipartite graph - like machines and jobs in the FFSP. Kwon et al. (2021) train MatNet using POMO (Kwon et al., 2020).
CPLEX
Besides the aforementioned MatNet model, we implement the mathematical model described above in the exact solver CPLEX with a time budget of 60 and 600 seconds per instance. However, with both time budgets, CPLEX is only capable of generating solutions to the FFSP20 instances.
Random and Shortest Job First
Third, we use different (meta-)heuristics to benchmark PARCO on the FFSP. The Random and Shortest Job First (SJF) heuristics are simple construction strategies that build valid schedules in an iterative manner. Starting from an empty schedule, the Random construction heuristic iterates through time steps $t=0,\dots T$ and stages $i=1\dots S$ and randomly assigns jobs available at the given time to an idle machine of the respective stage until all jobs are scheduled. Likewise, the SJF proceeds by assigning job-machine pairs with the shortest processing time first.
Genetic Algorithm
The Genetic Algorithms (GA) are metaheuristics widely used by the OR community to tackle the FFSP (Kahraman et al., 2008). The GA iteratively improves multiple candidate solutions called chromosomes. Each chromosome consists of $S\times N$ real numbers, where $S$ is the number of stages and $N$ is the number of jobs. For each job at each stage, the integer part of the corresponding number indicates the assigned machine index, while the fractional part determines job priority when multiple jobs are available simultaneously. Child chromosomes are created through crossover, inheriting integer and fractional parts independently from two parents. Mutations, applied with a 30% chance, use one of four randomly selected methods: exchange, inverse, insert, or change. The implementation uses 25 chromosomes. One initial chromosome is set to the Shortest Job First (SJF) heuristic solution and the best-performing chromosome is preserved across iterations. Each instance runs for 1,000 iterations.
Particle Swarm Optimization
Finally, Particle Swarm Optimization (PSO) iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration (Singh and Mahapatra, 2012). In this implementation, particles use the same representation as GA chromosomes. The algorithm employs 25 particles, with an inertial weight of 0.7 and cognitive and social constants set to 1.5. One initial particle represents the SJF heuristic solution. Like GA, PSO runs for 1,000 iterations per instance.
B.3.2 Datasets
Train data generation
We follow the instance generation scheme outlined in Kwon et al. (2021) sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$. For the main experiments on the FFSP shown in Table 3, we also use the same instance sizes as Kwon et al. (2021) with $N=20,50\text{ and }100$ jobs (denoted FFSP20, FFSP50 and FFSP100, respectively) that need to be processed in $S=3$ stages on $m_{i}=4$ machines each (i.e. $m=12$ machines in total). For the agent sensitivity analysis shown in Table 4, we fix the number of jobs to 50 but alter the number of agents. Still, we use $S=3$ for this experiment, but alter the number of machines per stage to $m_{i}=6,8\text{ and }10$, yielding a total of 18, 24 and 30 agents, respectively.
Testing
Testing is performed on 1,000 test instances generated according to the above sampling scheme. We use the same instances as Kwon et al. (2021) for the experiments of Table 3 and generate new ones for the experiments of Table 4.
B.3.3 PARCO Network Hyperparameters
Encoder
To solve the FFSP with our PARCO method, we use a similar encoder as Kwon et al. (2021). The MatNet encoder generates embeddings for all machines of all stages and the jobs they need to process, plus an additional dummy job embedding, which can be selected by any machine in each decoding step to skip to the next step. To compare PARCO with MatNet, we use similar hyperparameters for both models. We use $L=3$ encoder layers, generating embeddings of dimensionality $d_{h}=256$, which are split over $h=16$ attention heads in the MHA layers. Further, we employ Instance Normalization (Ulyanov et al., 2016) and a feed-forward network with 512 neurons in the transformer blocks of the encoder.
Decoder
The machines are regarded as the agents in our PARCO framework. As such, their embeddings are used as queries $Q_{t}^{s}$ in the Multiple Pointer Mechanism Eq. 5, while job embeddings are used as the keys and values. In each decoding step, the machine embeddings are fused with a projection of the time the respective machine becomes idle. Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Eq. 4.
Communication Layer
We employ a single transformer block with hidden dimension $d_{h}=256$ and $h=16$ attention heads in the MHA, an MLP with $512$ hidden units and Instance Normalization.
Agent Handler
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action is the highest. Formally, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
B.3.4 PARCO Training Hyperparameters
Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Eq. 9. We use the ADAM optimizer (Kingma and Ba, 2014) with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme. We train separate models for FFSP20, FFSP50 and FFSP100 instances for 100, 150 and 200 epochs, respectively, using 1,000 randomly generated instances per epoch split into batches of size 50.111Note: to avoid OOMs, For FFSP100 instances, batches are further split into mini-batches of size 25 whose gradients are accumulated. For the learned methods, the times given in parenthesis indicate the time required to train the respective models, whereas the other time information indicates the time required to solve a single instance. For the agent sensitivity analysis in Table 4, Matnet, as well as our PARCO model, are trained for 20 epochs on 1,000 instances per epoch. Duration in this table indicates the time required for training the respective models on the hardware specified in Section B.4.
B.3.5 Diagram for MatNet Decoding vs. PARCO Decoding for the FFSP
The following figures visualize the decoding for the machines of a given stage using MatNet and PARCO. As one can see in Fig. 6, MatNet requires a decoder forward pass for each machine to schedule a job on each of them. In contrast, as detailed in Fig. 6, PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
Figure 5: A FFSP Decoding Step with MatNet
Figure 6: A FFSP Decoding Step with PARCO
B.4 Hardware and Software
B.4.1 Hardware
Experiments were carried out on a machine equipped with two Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz CPUs with a total $128$ threads and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM. During inference, we employ only one CPU and a single GPU.
B.4.2 Software
We used Python 3.12, PyTorch 2.4 (Paszke et al., 2019) coupled with PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) with most code based on the RL4CO library (Berto et al., 2024a). The main OS is Ubuntu 22.04.4 LTS. There were no major difficulties when trying the code on other platforms, so we expect experiments to be reproducible in other architectures.
B.5 Source Code
We value open reproducibility. We provide the source code to reproduce our experiments on Github at https://github.com/ai4co/parco.
Appendix C Additional Discussion
Following Section 6, we outline further discussion here. We additionally remark that several areas of multi-agent planning, including multi-agent pathfinding (Lin and Ma, 2023; Wang et al., 2023; Tang et al., 2024a) and social robot navigation (Li et al., 2024a; He et al., 2024) employ communication as a tool with great importance, from which we also borrow ideas such as prioritized conflict resolution (Tang et al., 2024b).
An exciting future direction going forward is the development of foundation models, in which PARCO could aid in improving the efficiency and performance in solving multi-agent CO problems. Recent examples in the literature have demonstrated the possibility of scaling up solvers with (self)supervision (Drakulic et al., 2023; Luo et al., 2023; Corsini et al., 2024; Pirnay and Grimm, 2024; Luo et al., 2024), developing multi-task solvers for VRPs (Liu et al., 2024a; Lin et al., 2024; Zhou et al., 2024; Berto et al., 2024b) as well as more general CO tasks (Drakulic et al., 2024; Jiang et al., 2024). Such approaches model CO problems as sequential autoregressive decision-making; parallelizing this process with PARCO could enrich solution speed/quality. We note that most NCO approaches in e.g., VRPs, mostly focus on solving without explicitly modeling (heterogeneous) agents as in PARCO, which is of practical relevance in several real-world tasks. Better heuristics for prioritized planning might also be found utilizing automated discovery via recent LLM approaches for CO (Liu et al., 2024b; Ye et al., 2024b). Finally, integrating PARCO in online optimization (Hottung et al., 2021; Choo et al., 2022) and end-to-end pipelines for construction and improvement, as done recently in (Berto et al., 2024a; Kong et al., 2024), would further improve the performance of PARCO given larger computational time budgets.